Hi I need help on these questions.
"Determine the relationships and express them using variables."
Input--Output
12--16
16--19
20--22
24--25
26--26.5
Input--Output
1--3
2--5
3--9
4--17
5--33

Answer by stanbon(75887)

"Determine the relationships and express them using variables."
Input--Output
12--16
16--19
20--22
24--25
26--26.5
------
Notice that Output increases by 3 when Input increases by 4
So slope = 3/4
O = (3/4)I + b
Using (12,16), solve for "b""
16 = (3/4)12 + b
b = 16-9 = 7
Equation:: O = (3/4)I + 7
---------------------------------
Input--Output
1--3
2--5
3--9
4--17
5--33
----
Notice that output increases by 2 when input increases by 1.
slope = 2/1 = 2
----
If input is zero, output would be 1
----
Ans: O = 2I + 1
-------------------
Cheers,
Stan H.
--------------